X³/2 +2/x³=1 find x³.x³.x³ i get ans8 plz solve this

Grade:12th pass

2 Answers

subrat
6 years ago
First of all x can`t be real as can be seen by AM-GM inequality. LHS must >=2 or
Lucky
6 years ago
Let x^3/2 =t.Therefore, t+1/t =1 .t^2 +1 =t. -`a`So, x^3 . x^3. x^3 = 8t (t = x^3/2)= 8 ( t^2 +1) (from above)So t^2+1 is greater than or equal to `0` as it is positive. So least value will occur when t=0. So when we put t=0 in above we get x^3.x^3.x^3 =8(t^2 +1) = 8 is the required answer.